niahmk:

The net of an isosceles triangular prism is shown. What is the surface area, in square units, of the triangular prism? Three rectangles are shown, one below the other. The middle rectangle has a width of 6 units, and, along each width, is a triangle with a height of 4 units. The length of the lowermost rectangle is 8 units and its width is 5 units. 104 units2 132 units2 152 units2 168 units2

niahmk:

1 attachment
snowflake0531:

the same colored ones means that they're equal

1 attachment
snowflake0531:

let's first find the area of the purple one, what is the length and width?

niahmk:

6 and 6?

snowflake0531:

ummm relook one side is 6 yes, but the other side, relook

niahmk:

8

snowflake0531:

Yes, so what's the area of that rectangle

niahmk:

48

snowflake0531:

so now what's the area of the red rectangles

1 attachment
niahmk:

8x5?

snowflake0531:

yep so what's the area?

niahmk:

40

snowflake0531:

yes and the area of a triangle is base times height diivded by 2 What is the area of the blue triangle

1 attachment
niahmk:

4x4?

snowflake0531:

mmm the height is 4 yes, relook at the base, maybe look at the right blue triangle

niahmk:

the base is 6?

snowflake0531:

Yes so what's the area

niahmk:

4x6= 24

snowflake0531:

yes then divide that by half

snowflake0531:

Because area of a triangle is base time height then divided by 2

niahmk:

so 24 divided by 2?

snowflake0531:

yes

niahmk:

12

snowflake0531:

1 attachment
snowflake0531:

So now add them up 48+40+40+12+12

niahmk:

40+40+48+12+12= 52

niahmk:

so its C, thank youu

snowflake0531:

@niahmk wrote:
40+40+48+12+12= 52
*152 but yes, C

snowflake0531:

yw~